What is the truth value of ∀xP(x), where P(x) is the statement “x2 < 10” and the domain consists

of the positive integers not exceeding 4?



1
Expert's answer
2021-12-06T16:05:36-0500

"x^2<10\implies x\in (-\sqrt{10},\sqrt{10})". "4>\sqrt{10}" , so, P(4) = 0, so, the given statement is false